[Hallc_running] [New Logentry] Hall C RC update: Thursday, 7/18

gaskelld at jlab.org gaskelld at jlab.org
Thu Jul 18 10:20:03 EDT 2019

Logentry Text:
This morning, we completed taking data for the Q2=0.375, high epsilon setting. We took a Controlled Access during Hall A PQB work and rotated the spectrometers to their initial angles for the Q2=0.425 high epsilon setting (neither spectrometer is at their minimum angle, but the initial setting has the separation between the 2 at their smallest value - 18.6 degrees). There are no special tasks planned for today - we will take data at Q2=0.425. Upon completion of that kinematic setting, we will take some H(e,e'p) data, then move on the Q2=2.12 Pion Scaling setting. 

There will be an RC meeting today at 4:30 pm in the meeting room in the Counting House.  Bluejeans link: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3706233
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190718/b6e9715d/attachment.html>

More information about the Hallc_running mailing list